OpenStudy (anonymous):

Granite College has 1,200 female students. This represents 40 percent of the total student body. The number of male students at Granite is

OpenStudy (anonymous):

female=40% males=60% 40 * 3/2 =60 female * 3/2=male 1200 * 3/2 =1800